OpenStudy (anonymous):

(square root 2- square root 3)^2

OpenStudy (anonymous):

\[=5-2\sqrt{6}\]

OpenStudy (anonymous):

howd you get that

OpenStudy (anonymous):

\[(\sqrt{2}-\sqrt{3})(\sqrt{2}-\sqrt{3})=(\sqrt{2}-\sqrt{3})\sqrt{2}-(\sqrt{2}-\sqrt{3})\sqrt{3}=\sqrt{2}\sqrt{2}-\sqrt{3}\sqrt{2}-\sqrt{2}\sqrt{3}+\sqrt{3}\sqrt{3}=2-\sqrt{2}\sqrt{3}-\sqrt{2}\sqrt{3}+3=5-2\sqrt{2}\sqrt{3}=5-2\sqrt{6}\]

OpenStudy (anonymous):

so after that step you get -5-2square root 6?

OpenStudy (anonymous):

No positive five, because two negatives makes a positive

OpenStudy (anonymous):

okay iget

OpenStudy (anonymous):

thanks